# CMind: An AI Agent for Localizing C Memory Bugs

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 34th IEEE/ACM International Conference on Program Comprehension; April 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Bio-inspired approachesccs: Software and its engineering Software libraries and repositories
@@ 1. Introduction
We make CMind available as a command-line Unix tool and via a web platform. By default CMind uses GPT-o4 for LLM interaction but this is configurable and in principle any LLM can be used. For the web platform, the user needs only to upload an archive snapshot of his or her code and the text of a bug report. For the command-line tool, the user needs only to point CMind to the source code folder (or archive file) and a text file containing the bug report. In a pilot evaluation we demonstrate how to use the tool on twenty actual memory bugs in C programs we found from public repositories. CMind is open-source and available via:
@@ 2. Background/Related Work
Spectrum-based methods use run-time behaviors to localize bugs. For example, Jones et al. (Jones and Harrold, 2005) proposed Tarantula that uses the coverage of the test data for bug localization. Model-based approaches use probabilistic models. Baah et al. uses probabilistic graphical models to find bugs in the program. Information retrieval-based methods use static information for bug localization. Mahmud et al. (Mahmud et al., 2024) proposed an information retrieval approach that retrieves files with bug based on GUI information. Zhou et al. (Zhou et al., 2012) used bug reports to find the buggy code. Neural-based methods mean localize the bugs with neural network. Qian et al. (Qian et al., 2023) used graph convolution neural network for bug localization. Bio-based methods use clues from human behavior for bug localization such as the work by Muller et al. (Müller and Fritz, 2016) that uses bio-metrics for bug localization. More recently, several LLMs and agentic-based approaches have been proposed. Stein et al. (Stein et al., 2025) used LLMs for bug localization by exploring the attention in LLMs. Chang et al. (Chang et al., 2025) proposed LLM-based hierarchical framework for bug localization. Qin et al. (Qin et al., 2024) proposed a multi-agent based method for bug localization. Due to page limit, several other papers along these lines are highlighted in different survey papers (Niu et al., 2025; Wong et al., 2016).
Although LLMs have shown success in bug localization (Stein et al., 2025; Dai et al., 2025; Chang et al., 2025; Liu et al., 2024; Qin et al., 2024), LLMs still struggle with hallucinations (Li et al., 2025; Zhang et al., 2025) and understanding long contexts (Kuratov et al., 2024; Li et al., 2024). To limit the range of possible answers, we developed our method based on a human study with human programmers (Smith et al., 2025). Compared with other agentic based methods such as (Qin et al., 2024), the key novelty of this paper is that we design our prompt based on a human study with human programmers.
@@ 3.1. Information Collection
Area 1 in Figure 1 collects the necessary information. The input is the bug report and the source code. The output is the related functions or source files. First, we use LLMs to find the program entry points based on the bug report. The entry point that LLMs output could be functions names or files. We limit LLMs to find up to three related functions or files (including the entry function/file) at this time. We limit the output to three related functions and files because Zheng et al. (Zheng et al., 2025) observed that LLMs are prone to hallucinate in free-form and longer outputs. Then, we used the Python script to obtain all the functions and files requested by LLMs. Note that we do not provide LLMs the entire project because of the context limit although LLMs can have large context size. We used the following prompt form for information collection LLMs, where {bug_report} is the bug report and {filename} is filenames of source code files.
@@ 3.2. Static Analysis
Static analysis is as shown in Area 2 in Figure 1. We first used LLMs to decide which static analysis tasks to use. Then, we used tools to generate static analysis results based on the decision that LLMs make. We used tools to generate the static analysis results because LLMs are susceptible to hallucinate on static analysis tasks (Su and McMillan, 2025) and more complex source code (Xie et al., 2025). LLMs are only responsible for reasoning which static analysis task to use and which paths to look at. We limited LLMs to two choices, i.e. callgraph generation and dataflow analysis. We limited LLMs to only two choices because these are two main strategies that human programmers use for bug localization (Ko and Myers, 2004; Smith et al., 2025). In addition, LLMs are required to provide the source and sink when dataflow analysis is chosen. We used Joern (joern, 2025) for dataflow analysis and Doxygen (van Heesch, 2013) for callgraph generation. We used Joern and Doxygen because these tools are used in several research papers (Steenhoek et al., 2024; Chen et al., 2023; CSÁSZÁR and SLAVESCU, 2020). To reduce the length of the tokens, we used the same LLMs to find the most related call chains in the callgraph. This is because the size of callgraph could potentially become very large and not all of the call chains are related. We use the following prompt for deciding which static analysis to use, where {bug_report} is the bug report from the user, {codeblocks} is the source code of the entry point, and {function_name} is the name of the entry point functions.
Now you can use call graph analysis, a method to analyze how the methods call each other and data flow analysis, a method to analyze how data flow in the problem. Given the bug report {bug_report} and the related methods {codeblocks}, could you tell me which method would you use to analyze the bugs? Please just tell me what you would use and be sure to refer to the context and not make up anything yourself. Please use the template: data flow analysis: source:[SOURCE] sink:[SINk] if you think data flow analysis is an appropriate method. Please note that source should be the name of the method in {function_name} and sink should be in the a function that is inside the same method (if this is not the case, please suggest call graph) and do not provide any other information and the example is data flow analysis: source:[A] sink:[B] where A is the method name and B is the function name. otherwise: call graph analysis
@@ 6. Future Works
We divide our future works into three directions. First, we aim to integrate this tool into IDE/editors such as VS code for better usability. Second, we aim to increase the robustness of our approach. Currently, CMind could only localize bugs with a clear stack trace. We aim to improve CMind by conducting experiments with bug reports that have more implicit symptoms with thorough analysis and bug categorization. Third, we aim to conduct the experiments with more diverse models (both open-source models and close-models) to improve the generalizability of prompts.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 34th IEEE/ACM International Conference on Program Comprehension; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: 34th IEEE/ACM International Conference on Program Comprehension (ICPC ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3794763.3794829isbn: 979-8-4007-2482-4/2026/04ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software notations and tools
@@ 1. Introduction
We make CMind available as a command-line Unix tool and via a web platform. By default CMind uses GPT-o4 for LLM interaction but this is configurable and in principle any LLM can be used. For the web platform, the user needs only to upload an archive snapshot of his or her code and the text of a bug report. For the command-line tool, the user needs only to point CMind to the source code folder (or archive file) and a text file containing the bug report. In a pilot evaluation we demonstrate how to use the tool on twenty actual memory bugs in C programs we found from public repositories.
@@ 2. Background/Related Work
Spectrum-based methods use run-time behaviors to localize bugs. For example, Jones et al. (Jones and Harrold, 2005) proposed Tarantula that uses the coverage of the test data for bug localization. Model-based approaches use probabilistic models. Baah et al. uses probabilistic graphical models to find bugs in the program. Information retrieval-based methods use static information for bug localization. Mahmud et al. (Mahmud et al., 2024) proposed an information retrieval approach that retrieves files with bug based on GUI information. Zhou et al. (Zhou et al., 2012) used bug reports to find the buggy code. Neural-based methods mean localizing the bugs with neural network. Qian et al. (Qian et al., 2023) used graph convolution neural network for bug localization. Bio-based methods use clues from human behavior for bug localization such as the work by Muller et al. (Müller and Fritz, 2016) that uses bio-metrics for bug localization. More recently, several LLMs and agentic-based approaches have been proposed. Stein et al. (Stein et al., 2025) used LLMs for bug localization by exploring the attention in LLMs. Chang et al. (Chang et al., 2025) proposed LLM-based hierarchical framework for bug localization. Due to page limit, several other papers along these lines are highlighted in different survey papers (Niu et al., 2025; Wong et al., 2016).
Although LLMs have shown success in bug localization (Stein et al., 2025; Dai et al., 2025; Chang et al., 2025; Liu et al., 2024; Qin et al., 2024), LLMs still struggle with hallucinations (Li et al., 2025; Zhang et al., 2025) and understanding long contexts (Kuratov et al., 2024; Li et al., 2024). To limit the range of possible answers, we developed our method based on a human study with human programmers (Smith et al., 2025). Specifically, we extracted the information that programmers used with tools during the study and limited LLMs to only look at the information used by the programmers. Compared with other agentic-based methods such as (Qin et al., 2024), the key novelty of this paper is that we design our prompt based on a human study with human programmers.
@@ 3.1. Information Collection
Area 1 in Figure 1 collects the necessary information. The input is the bug report and the source code. The output is the related functions or source files. We use LLMs to find the program entry points based on the bug report. The entry point that LLMs output could be functions names or files related to the bug report. We limit LLMs to find up to three related functions or files (including the entry function/file) at this time. We limit the output to three related functions and files because Zheng et al. (Zheng et al., 2025) observed that LLMs are prone to hallucinate in free-form and longer outputs. We used the Python script to obtain all the functions and files requested by LLMs. Note that we do not provide LLMs the entire project because of the context size limit of LLMs although LLMs can have large context size. We used the following prompt form for information collection LLMs, where {bug_report} is the bug report and {filename} is filenames of source code files.
@@ 3.2. Static Analysis
Static analysis is as shown in Area 2 in Figure 1. We first used LLMs to decide which static analysis tasks to use. Then, we used tools to generate static analysis results based on the decision that LLMs make. We used tools to generate the static analysis results because LLMs are susceptible to hallucinate on static analysis tasks (Su and McMillan, 2025) and more complex source code (Xie et al., 2025). LLMs are only responsible for reasoning which static analysis task to use and which paths to look at. We limited LLMs to two choices, i.e., callgraph generation and dataflow analysis. We limited LLMs to only two choices because these are two main strategies that human programmers use for bug localization (Ko and Myers, 2004; Smith et al., 2025). In addition, LLMs are required to provide the source and sink when dataflow analysis is chosen. We used Joern (joern, 2025) for dataflow analysis and Doxygen (van Heesch, 2013) for callgraph generation. We used Joern and Doxygen because these tools are used in several research papers (Steenhoek et al., 2024; Chen et al., 2023; CSÁSZÁR and SLAVESCU, 2020). To reduce the length of the tokens, we used the same LLMs to find the most related call chains in the callgraph. This is because the size of callgraph could potentially become very large and not all of the call chains are related. We use the following prompt for deciding which static analysis to use, where {bug_report} is the bug report from the user, {codeblocks} is the source code of the entry point, and {function_name} is the name of the entry point functions.
Now, you can use call graph analysis, a method to analyze how the methods call each other and data flow analysis, a method to analyze how data flow in the problem. Given the bug report {bug_report} and the related methods {codeblocks}, could you tell me which method would you use to analyze the bugs? Please just tell me what you would use and be sure to refer to the context and not make up anything yourself. Please use the template: data flow analysis: source:[SOURCE] sink:[SINK] if you think data flow analysis is an appropriate method. Please note that source should be the name of the method in {function_name} and sink should be in a function that is inside the same method (if this is not the case, please suggest call graph) and do not provide any other information and the example is data flow analysis: source:[A] sink:[B] where A is the method name and B is the function name. otherwise: call graph analysis
@@ 6. Future Works
We divide our future works into four directions. First, we aim to integrate CMind into IDE/editors such as VS code to increase the usefulness of this tool. Second, we aim to increase the robustness of our approach. Currently, CMind could only localize bugs with clear stack trace. We aim to improve CMind by conducting experiments with more diverse bug reports and complete analysis, so it could localize bugs with implicit symptoms. Third, we aim to conduct the experiments with more models to improve the generalizability of our prompts. Finally, we aim to develop a model that can be run locally to protect the data privacy.